OpenStudy (anonymous):

abs(6-3i) it is 3squareroot5 but why? abs(6-3i) it is 3squareroot5 but why? @Mathematics

OpenStudy (anonymous):

\[\left| 6-3i \right| is 3\sqrt{5}\]

myininaya (myininaya):

\[\sqrt{6^2+3^2}=\sqrt{36+9}=\sqrt{45}=\sqrt{9 \cdot 5} =3 \sqrt{5}\]

myininaya (myininaya):

\[|x+yi|=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

oh ok thanks